# RoSIS: Robust Framework for Text-Promptable Surgical Instrument Segmentation Using Vision-Language Fusion

### v1

@@ 4.3.2 Comparison on the EndoVis2017
Table 4: Visual and language feature fusion structural ablation study on EndoVis2018 dataset.
ISI IoU mc IoU
VLTF [15] 68.58 33.48
PWAM [12] 66.94 36.89
Ours 71.65 38.23
(+) Raw language feature 74.36 (+2.71) 41.87 (+3.64)
(+) Spatial token 66.76 (-4.89) 27.35 (-10.88)
(+) Language token 76.16 (+4.51) 44.54 (+6.31)
@@
4.3.3 Ablation study of the prompt and iterative inference process
Table 3 presents the ablation study results on the EndoVis2018 dataset, emphasizing the role of iterative inference and text prompts. Starting with the baseline using only the name prompt such as “The bipolar forceps,” the model achieves ISI IoU and mc IoU scores of 67.74 and 37.84, respectively. Adding the location prompt to the name prompt slightly improves performance in ISI IoU. Then, the iterative inference consistently improves performance across both metrics. When applied to the location prompt configuration, iterative inference significantly boosts ISI IoU and mc IoU. This demonstrates that iterative inference effectively refines the segmentation map by utilizing location-specific prompts, contributing to more accurate segmentation. Similarly, when iterative inference is incorporated with the gpt4 prompt, which provides descriptive characteristics for each instrument, there is a substantial improvement in segmentation accuracy. The ISI IoU increases by 4.79, and mc IoU increases by 3.93. This improvement shows the synergy between descriptive prompts and iterative inference, as the iterative process allows the model to refine segmentation based on more detailed object information. Also, as shown in the red circle part of Fig. 5, iterative inference allows for more accurate object segmentation maps.
These results indicate the crucial role of iterative inference in our model. By refining segmentation through multiple stages and adapting prompts based on previous outputs, iterative inference contributes to significant gains in both general and class-specific accuracy, making it a key factor in achieving optimal segmentation performance.
4.3.4 Ablation study of the structure
Table 4 presents the results of our ablation study on the EndoVis2018 dataset, evaluating the impact of different fusion structures on the integration of visual and language features. We use ISI IoU and mc IoU as comparison metrics to focus on performance in robust settings for all ablation studies. The baseline models, VLTF [15] and PWAM [12], achieve ISI IoU scores of 68.58 and 66.94, and mc IoU scores of 33.48 and 36.89, respectively. In comparison, our model with the base fusion structure performs better, achieving an ISI IoU of 71.65 and an mc IoU of 38.23.
Further modifications demonstrate the effectiveness of specific fusion components. Adding raw language features increases ISI IoU and mc IoU to +2.71 and +3.64, respectively. The raw language feature means using the language feature from BERT [29] directly as input for the MHCA in the existence prediction branch. Additionally, incorporating spatial tokens reduces ISI IoU by -4.89 and mc IoU by -10.88, suggesting that spatial tokens may introduce irrelevant spatial information, which negatively impacts performance. A spatial token refers to a structure that applies MHCA to visual features instead of language tokens in the MMFB. These results show that it is important to retain the visual features that have the original spatial and structural information from the visual encoder when performing vision-language feature fusion. Conversely, adding language tokens significantly improves results, with ISI IoU reaching 76.16 (+4.51) and mc IoU reaching 44.54 (+6.31), confirming that structured language information enhances the model’s ability to segment surgical instruments accurately.
These findings demonstrate the importance of selecting the appropriate fusion strategies. Language tokens are the most effective, providing a robust integration of language and visual features that significantly enhance segmentation performance.
4.3.5 Qualitative results
@@
In Fig. 6, we visualize the prediction of our model, the groud truth, and the other promptable segmentation methods [12, 10]. Since other methods are trained without robust conditions, they hard to determine the presence or absence of objects in the R-SIS environment, resulting in higher false positives in the predicted maps. As shown in the 4th and 5th columns of Fig. 6, the segmentation masks for various classes overlap. This overlap occurs because segmentation maps are generated for all classes’ text prompts, causing all maps to merge. In contrast, our model first verifies the presence or absence of an object before generating a map, producing a clean segmentation for a single object and leading to a low false positive rate, which is crucial in surgical environments.
4.3.6 Cross-dataset study
We also evaluate our method in a cross-dataset setting. In this setup, the model is trained on the EndoVis2018 dataset and tested on the EndoVis2017 dataset, and vice versa. The EndoVis2017 and EndoVis2018 datasets share five common categories and each includes two unique categories. As shown in Tables 5 and 6, RoSIS demonstrates better performance compared to other promptable segmentation methods. Additionally, even among shared categories, there are slight differences in shape and domain variations, making the cross-dataset experiment indicate the generalizability of methods.
The promptable segmentation methods LAVT [12] and TP-SIS [10] show overall lower performance, with a large discrepancy between their Ch IoU and ISI IoU scores. In contrast, our method, RoSIS, achieves the best results, with a smaller gap between Ch IoU and ISI IoU, highlighting its robustness and efficiency. These results demonstrate the generalizability and effectiveness of RoSIS in handling domain shifts across different datasets.
@@ 4.3.6 Cross-dataset study
Table 5: Cross-dataset setting performance comparison by training on the EndoVis2018 and testing on the EndoVis2017.
Method Ch IoU ISI IoU mc IoU
LAVT [12] 40.82 14.35 13.26
@@
Table 6: Cross-dataset setting performance comparison by training on the EndoVis2017 and testing on the EndoVis2018.
Method Ch IoU ISI IoU mc IoU
LAVT [12] 46.57 15.67 14.91
@@ 5 Conclusion
In this work, we presented Robust Surgical Instrument Segmentation (RoSIS), a novel framework focused on enhancing the robustness and effectiveness of text-promptable surgical instrument segmentation. Recognizing the limitations of existing methods, which often rely on assumptions about object presence, we redefined the task under robust conditions as Robust text-promptable SIS (R-SIS). RoSIS employs an encoder-decoder architecture with a Multi-Modal Fusion Block (MMFB) and a Selective Gate Block (SGB) to achieve a balanced integration of visual and language features, preventing the dominance of either modality. Additionally, we introduced an iterative inference strategy to refine segmentation maps, utilizing both name-based and location-based prompts to improve accuracy. Our experiments on the EndoVis2017 and EndoVis2018 datasets demonstrated that RoSIS outperforms existing vision-based and promptable methods under robust conditions. Ablation studies further validated the effectiveness of our fusion structures, as well as the iterative inference approach, confirming RoSIS’s adaptability across diverse surgical scenarios. This work lays a foundation for future research in multi-modal surgical instrument segmentation, with applications in other areas where vision-language integration is essential for complex surgical scene understanding.

### v3

@@ 4.4.1 Iterative refinement process and prompt design
Table 7: A visual and language feature fusion structural ablation study and performance comparison of the fusion structure on the EndoVis2018 dataset.
Method MMFB SGB RL LT ISI IoU Ch IoU mc IoU
Ours 64.31 72.75 34.27
71.65 77.07 38.23
74.36 80.37 41.87
76.16 77.66 44.54
VLTF (RefSegformer) [16] 68.58 74.19 33.48
PWAM (LAVT) [20] 66.94 79.34 36.89
4.4.2 Structure design
To evaluate the effectiveness of our visual-language fusion structure, we conduct an ablation study on the EndoVis2018 dataset, analyzing the impact of each module on segmentation performance. Specifically, we assess the contributions of the Multi-Modal Fusion Block (MMFB), Selective Gate Block (SGB), Raw Language features (RL), and Language Token (LT) by incrementally adding components to our framework. RL involves using language features extracted from BERT [22] directly as input for the MHCA module in the existence prediction branch, enhancing textual representation within our segmentation framework. The results are summarized in Table 7.
Our findings demonstrate that each proposed module contributes to segmentation performance. SGB improves segmentation stability by selectively refining the fusion of visual and language features. RL enhances the model’s ability to process textual information, leading to slight improvements in segmentation accuracy. LT further strengthens feature representation, helping the model to refine segmentation masks more effectively. When combined, these modules achieve the best overall performance, validating their role in improving segmentation robustness.
To evaluate the effectiveness of our fusion structure, we compare our model with previous RIS methods, including VLTF (RefSegformer) [16] and PWAM (LAVT) [20]. The results show that our method improves segmentation accuracy and maintains stable performance. Unlike previous methods that rely heavily on either vision or language cues, our approach balances multi-modal information more effectively, reducing misclassification and false positives. The consistent improvement across key metrics confirms its robustness and reliability in surgical instrument segmentation.
@@ 4.4.2 Structure design
4.5 Qualitative results
@@
In Fig. 5, we visualize the prediction of our model, the groud truth, and the other promptable segmentation methods [20, 10]. Since other methods are trained without robust conditions, they hard to determine the presence or absence of instruments in the R-SIS environment, resulting in higher false positives in the predicted maps. As shown in the 4th and 5th columns of Fig. 5, the segmentation masks for various classes overlap. This overlap occurs because segmentation maps are generated for all classes’ text prompts, causing all maps to merge. In contrast, our model first verifies the presence or absence of an instrument before generating a map, producing a clean segmentation for a single instrument and leading to a low false positive rate, which is crucial in surgical environments.
@@
Table 8: Cross-dataset setting performance comparison by training on the EndoVis2018 and testing on the EndoVis2017, and vice versa.
Setting Method Ch IoU ISI IoU mc IoU
EV18$\rightarrow$ EV17 LAVT [20] 40.82 14.35 13.26
@@
[…]
@@
4.6 Cross-dataset study
We also evaluate our method in a cross-dataset setting. In this setup, the model is trained on the EndoVis2018 dataset and tested on the EndoVis2017 dataset, and vice versa. The EndoVis2017 and EndoVis2018 datasets share five common categories and each includes two unique categories. As shown in Table 8, RoSIS demonstrates better performance compared to other promptable segmentation methods. Additionally, even among shared categories, there are slight differences in shape and domain variations, making the cross-dataset experiment indicate the generalizability of methods.
The promptable segmentation methods LAVT [20] and TP-SIS [10] show overall lower performance, with a large discrepancy between their Ch IoU and ISI IoU scores. In contrast, our method, RoSIS, achieves the best results, with a smaller gap between Ch IoU and ISI IoU, showing its robustness and efficiency. These results demonstrate the generalizability and effectiveness of RoSIS in handling domain shifts across different datasets.
@@ 5 Conclusion
In this work, we redefined the task of text-promptable Surgical Instrument Segmentation (SIS) under robust conditions as Robust text-promptable SIS (R-SIS). Existing methods often assume the presence of instruments described by prompts, leading to challenges in realistic surgical scenarios. Such assumptions introduce biases in evaluation, as they leverage prior knowledge that is not available to vision-based models. R-SIS eliminates this dependency by enforcing prompt-based segmentation without prior knowledge of instrument presence. This ensures fairer comparisons between promptable and vision-based methods while improving robustness in real-world surgical applications.
To support this task setting, we developed Robust Surgical Instrument Segmentation (RoSIS), effective framework designed to address prompt uncertainty and instrument absence. Rather than relying on architectural complexity, RoSIS integrates prompt types with an existence prediction mechanism and an iterative refinement strategy, enabling reliable segmentation under R-SIS constraints.
Our experiments on the EndoVis2017 and EndoVis2018 datasets demonstrate that RoSIS consistently reduces false positive predictions while maintaining strong precision and recall. Ablation studies further validate that even without existence prediction, RoSIS performs competitively against existing models, underscoring the value of our task-aware formulation. This work highlights the importance of fair evaluation protocols for promptable segmentation and provides a foundation for future research in multi-modal surgical scene understanding.
6 Acknowledgements
This work was supported by the Technology Innovation Program (RS-2024-00443054, Development of a Supermicrosurgical Robot System for Sub-0.8mm Vessel Anastomosis through Human-Robot Autonomous Collaboration in Surgical Workflow Recognition) funded By the Ministry of Trade Industry & Energy (MOTIE, Korea).